OpenStudy (anonymous):

How do you solve the Pythagorean Theorem??? C^2=10.4^2+16.9^2???

OpenStudy (anonymous):

What is 10.4^2? What is 16.9^2? Add those together. Once you've done that, take the square root of that number so that way you have c=???. So 10.4^2+16.9^=108.16+285.61=393.77=c^2. So take the square root of both sides you have, c= \[\sqrt(393.77)\] = 19.84363878....

OpenStudy (anonymous):

btw, do you know what Pythagorean Theorem is used for?

OpenStudy (anonymous):

You asking me or Sunshine?

OpenStudy (anonymous):

sunshine